How can I find the slope for this problem:
x= 2-y
____
9
Its 2-y over 9 thanks

Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

put in form y=mx+b
9x=2-y
y+9x=2
y=2x-9
slope is 2

Answer by Alan3354(69443)

9x = 2-y
y = -9x + 2
slope = -9